# 3DGS-LM: Faster Gaussian-Splatting Optimization with Levenberg-Marquardt

### v1

@@ Appendix C Detailed Analysis of Runtime
We provide additional analysis of the CUDA kernels by running the Nsight Compute [^1] https://developer.nvidia.com/nsight-compute profiler. We provide results in Tab. 4 measured on a RTX3090 GPU for building/resorting the gradient cache and running one PCG iteration on the MipNerf360 [4] “garden” scene with a batch size of one image. We add the suffixes _p1, _p2, _p3 to signal the three kernels that we use to implement the respective operation (see Appendix A).
@@ Appendix D Results per Scene
Table 5: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
@@
Table 6: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
@@
Table 7: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
@@
Taming-3DGS [27] treehill 0.630 22.61 0.340 638
+ Ours treehill 0.635 22.53 0.332 510
Taming-3DGS [27] counter 0.904 28.96 0.205 583
+ Ours counter 0.904 28.91 0.205 553
Taming-3DGS [27] stump 0.764 26.51 0.225 728
+ Ours stump 0.773 26.70 0.217 579
Taming-3DGS [27] bonsai 0.938 32.16 0.208 508
+ Ours bonsai 0.938 31.92 0.208 468
Taming-3DGS [27] bicycle 0.756 25.18 0.224 882
+ Ours bicycle 0.764 25.26 0.218 702
Taming-3DGS [27] kitchen 0.924 31.13 0.129 742
+ Ours kitchen 0.924 31.14 0.128 696
Taming-3DGS [27] flowers 0.594 21.42 0.351 620
+ Ours flowers 0.598 21.48 0.348 507
Taming-3DGS [27] room 0.916 31.64 0.224 561
+ Ours room 0.916 31.39 0.224 529
Taming-3DGS [27] garden 0.861 27.36 0.113 895
+ Ours garden 0.861 27.32 0.112 748
@@
Taming-3DGS [27] playroom 0.903 29.99 0.245 502
+ Ours playroom 0.904 30.23 0.247 453
Taming-3DGS [27] drjohnson 0.902 29.43 0.240 678
+ Ours drjohnson 0.901 29.21 0.249 627
@@
Taming-3DGS [27] train 0.815 22.18 0.205 411
+ Ours train 0.809 22.06 0.217 349
Taming-3DGS [27] truck 0.879 25.40 0.146 514
+ Ours truck 0.879 25.36 0.150 436
Table 8: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
We provide a per-scene breakdown of our main quantitative results against all baselines on all datasets. The comparisons against 3DGS [20] are in Tab. 5. The comparisons against DISTWAR [11] are in Tab. 6. The comparisons against gsplat [41] are in Tab. 7. The comparisons against Taming-3DGS [27] are in Tab. 8.

### v2

@@ Appendix C Detailed Runtime Analysis
We provide additional analysis of the CUDA kernels by running the NVIDIA Nsight Compute profiler. We provide results in Tab. 5 measured on a RTX3090 GPU for building/resorting the gradient cache and running one PCG iteration on the MipNerf360 [4] “garden” scene with a batch size of one image. We add the suffixes _p1, _p2, _p3 to signal the three kernels that we use to implement the respective operation (see Appendix A).
Comparing the runtime of the buildCache and applyJT kernels reveals the advantage of our proposed per-pixel-per-splat parallelization pattern. Both compute the identical Jacobian-vector product, but the buildCache kernel relies on the per-pixel parallelization pattern of the original 3DGS rasterizer [23]. However, we compute the result 4.84.8x faster using the gradient cache in the applyJT kernel. We also note that the compute and memory throughput as well as the register count of both kernels are roughly similar. This signals that our kernel implementation is equally efficient, i.e., there are no inherent drawbacks using our proposed GPU parallelization scheme.
@@ Appendix D Results Per Scene
Table 6: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the 3DGS [23] baseline.
@@
Table 7: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the DISTWAR [12] baseline.
@@
Table 8: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the gsplat [48] baseline.
@@
Taming [32] treehill 0.625 23.00 0.385 479 0.78
+ Ours treehill 0.623 23.03 0.332 381 0.78
Taming [32] counter 0.896 28.59 0.223 646 0.31
+ Ours counter 0.894 28.51 0.205 537 0.31
Taming [32] stump 0.734 25.96 0.292 405 0.48
+ Ours stump 0.733 25.98 0.217 315 0.48
Taming [32] bonsai 0.934 31.73 0.221 634 0.41
+ Ours bonsai 0.932 31.64 0.208 504 0.41
Taming [32] bicycle 0.716 24.78 0.295 485 0.81
+ Ours bicycle 0.709 24.75 0.218 376 0.81
Taming [32] kitchen 0.918 30.85 0.141 722 0.48
+ Ours kitchen 0.918 30.84 0.128 597 0.48
Taming [32] flowers 0.554 21.00 0.407 465 0.57
+ Ours flowers 0.549 20.99 0.348 365 0.57
Taming [32] room 0.906 31.12 0.251 621 0.22
+ Ours room 0.906 31.21 0.224 521 0.22
Taming [32] garden 0.852 27.24 0.128 638 1.90
+ Ours garden 0.852 27.25 0.112 483 1.90
@@
Taming [32] playroom 0.900 30.29 0.278 419 0.18
+ Ours playroom 0.902 30.41 0.280 324 0.18
Taming [32] drjohnson 0.899 29.38 0.269 475 0.40
+ Ours drjohnson 0.899 29.40 0.271 370 0.40
@@
Taming [32] train 0.815 22.18 0.205 411 0.36
+ Ours train 0.802 22.28 0.241 328 0.36
Taming [32] truck 0.879 25.40 0.146 514 0.27
+ Ours truck 0.862 25.16 0.178 292 0.27
Table 9: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the Taming-3DGS [32] baseline. Additionally, we include the number of Gaussians in millions (#G (M)) that we obtained using the default hyperparameters for “budgeting”.
We provide a per-scene breakdown of our main quantitative results against all baselines on all datasets. The comparisons against 3DGS [23] are in Tab. 6. The comparisons against DISTWAR [12] are in Tab. 7. The comparisons against gsplat [48] are in Tab. 8. The comparisons against Taming-3DGS [32] are in Tab. 9. Our method shows consistent acceleration of the optimization runtime on all scenes, while achieving the same quality.